OpenStudy (anonymous):

On Line DE D = (-3,-3) and E (2,4). Find point F on line DE so that DF and EF form a 1:3 ratio.

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

1:3 ratio means there are 4 segments on the line (1 + 3 = 4). Start by finding the "length" of each segment in x an y. \[x_{length}=\frac{ |x _2-x_1| }{ 4 }\] \[y_{length}=\frac{ |y _2-y_1| }{ 4 }\]

OpenStudy (anonymous):

xlength 5/4 and ylength= 7/5

OpenStudy (anonymous):

7/4 for y.

OpenStudy (anonymous):

soorry i meant 7/4

OpenStudy (anonymous):

and then since you want a 1:3 ratio, add those to point D one time to get F.